# Energy, enstrophy and helicity transfers in polymeric turbulence

### v1

@@ 2.1 The governing equations
$\begin{gathered}\frac{\partial u_{i}}{\partial t}+u_{j}\frac{\partial u_{i}}{\partial x_{j}}=-\frac{1}{\rho}\frac{\partial p}{\partial x_{i}}+\nu\frac{\partial^{2}u_{i}}{\partial x_{j}\partial x_{j}}+\frac{1}{\rho}\frac{\partial T_{ij}}{\partial x_{j}}+f_{i},\\ \frac{\partial R_{ij}}{\partial t}+u_{k}\frac{\partial R_{ij}}{\partial x_{k}}=\frac{\partial u_{i}}{\partial x_{k}}R_{kj}+R_{ik}\frac{\partial u_{k}}{\partial x_{j}}-\frac{1}{\tau_{p}}\left(\mathcal{P}R_{ij}-\delta_{ij}\right),\\ \frac{\partial u_{j}}{\partial x_{j}}=0,\end{gathered}$ (1)
@@
$R_{ij}=T_{ij}\frac{\tau_{p}}{\mu_{p}}+\delta_{ij},$ (2)
where $\tau_{p}$ is the polymeric relaxation time, $\mu_{p}$ is the polymeric viscosity and $\delta_{ij}$ is the Kronecker delta. $\mathcal{P}$ is the Peterlin function and is equal to $\mathcal{P}=1$ for the purely elastic Oldroyd-B model and to $\mathcal{P}=(\ell_{max}^{2}-3)/(\ell_{max}^{2}-R_{ii})$ for the FENE-P model; $\ell_{max}$ is the maximum polymer extensibility, $R_{ii}$ is the instantaneous measure of the end-to-end length of the polymers.
@@
$\frac{\partial\omega_{i}}{\partial t}+u_{j}\frac{\partial\omega_{i}}{\partial x_{j}}=\omega_{j}\frac{\partial u_{i}}{\partial x_{j}}-\nu\frac{\partial^{2}\omega_{i}}{\partial x_{j}\partial x_{j}}+\frac{1}{\rho}\frac{\partial T_{ij}^{\omega}}{\partial x_{j}}+f_{i}^{\omega},$ (3)
@@
[…]
@@ 2.2 The budget equations
Following the work of Baj et al. (2022), we consider three specific structure functions, i.e. the velocity structure function $\delta q^{2}=\delta u_{i}\delta u_{i}$, the vorticity structure function $\delta\omega^{2}=\delta\omega_{i}\delta\omega_{i}$ and the helicity structure function $\delta h=\delta u_{i}\delta\omega_{i}$ where repeated indices imply summation. The three structure functions feature velocity and vorticity incremements ($\delta u_{i}$ and $\delta\omega_{i}$) between two points $\bm{x}$ and $\bm{x}^{\prime}$, that can be identified by means of their midpoint $\bm{X}=(\bm{x}+\bm{x}^{\prime})/2$ and separation vector $\bm{r}=\bm{x}^{\prime}-\bm{x}$, namely
@@
The budget equations for $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ describe production, transport and dissipation of energy, enstrophy and helicity in the space of scales $\bm{r}$ and positions $\bm{X}$. These equations link the variation in time of the three structure functions at a given scale and position to the instantaneous unbalance among production, transport and dissipation. The three budget equations are obtained after manipulation of the Navier–Stokes equations for the velocity and the vorticity, without any assumption of homogeneity and isotropy; for the full derivation (for a purely Newtonian fluid) of the equation for $\left\langle{\delta q^{2}}\right\rangle$ we refer the interested reader to the appendix of Gatti et al. (2020). The derivation of the budget equations for $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ starts with the evolution equations for velocity and vorticity, and requires a sequence of manipulations that use the incompressibility constraint. A brief recap of the main steps is also provided in Baj et al. (2022).
@@
In compact form, the evolution equations for $\left\langle{\delta q^{2}}\right\rangle=\left\langle{\delta u_{i}\delta u_{i}}\right\rangle$, $\left\langle{\delta h}\right\rangle=\left\langle{\delta u_{i}\delta\omega_{i}}\right\rangle$ and $\left\langle{\delta\omega^{2}}\right\rangle=\left\langle{\delta\omega_{i}\delta\omega_{i}}\right\rangle$ can be written as:
@@
$\frac{\partial\left\langle{\delta g}\right\rangle}{\partial t}+\frac{\partial\psi^{\delta g}_{j}}{\partial X_{j}}+\frac{\partial\phi^{\delta g}_{j}}{\partial r_{j}}=P^{\delta g}+\Pi^{\delta g}+\mathcal{E}^{\delta g}+F^{\delta g},$ (4)
where $\delta g$ is a generic quantity. On the left hand side, $\bm{\psi}^{\delta g}$ and $\bm{\phi}^{\delta g}$ are the components of the six-dimensional vector field of fluxes $\bm{\Phi}^{\delta g}$ in the physical space $\bm{X}$ and in the space of scales $\bm{r}$ respectively. On the right hand side, $P^{\delta g}$ is the production term, $\Pi^{\delta g}$ denotes the exchange of $\delta g$ between the fluid phase and the polymeric microstructure at position $\bm{X}$ and scale $\bm{r}$, $\mathcal{E}^{\delta g}$ is the fluid dissipation term, and $F^{\delta g}$ denote the energy entering the system at a certain position and scale due to the external forcing.
@@
In the following, we detail the terms of the budget equation for the three structure functions. The components of the flux in the physical space $\bm{\psi}$ read
$\psi_{j}^{\delta q^{2}}=$ $\underbrace{\left\langle{U_{j}^{*}\delta q^{2}}\right\rangle}_{\text{Mean transp.}}+\underbrace{\left\langle{u_{j}^{*}\delta q^{2}}\right\rangle}_{\text{Turbulent transp.}}+\underbrace{\frac{2}{\rho}\left\langle{\delta p\delta u_{j}}\right\rangle}_{\text{Pressure transp.}}$ (5)
$\underbrace{-\frac{\nu}{2}\frac{\partial\left\langle{\delta q^{2}}\right\rangle}{\partial X_{j}}}_{\text{Viscous diff.}}+\underbrace{\frac{1}{\rho}\left\langle{\delta T_{ij}\delta u_{i}}\right\rangle,}_{\text{Polymeric transp.}}$
$\psi_{j}^{\delta\omega^{2}}=\underbrace{\left\langle{U_{j}^{*}\delta\omega^{2}}\right\rangle}_{\text{Mean transp.}}+\underbrace{\left\langle{u_{j}^{*}\delta\omega^{2}}\right\rangle}_{\text{Turbulent transp.}}\underbrace{-\frac{\nu}{2}\frac{\partial\left\langle{\delta\omega^{2}}\right\rangle}{\partial X_{j}}}_{\text{Viscous diff.}}+\underbrace{\frac{1}{\rho}\left\langle{\delta T^{\omega}_{ij}\delta\omega_{i}}\right\rangle,}_{\text{Polymeric transp.}}$ (6)
$\psi_{j}^{\delta h}=$ $\underbrace{\left\langle{U_{j}^{*}\delta h}\right\rangle}_{\text{Mean transp.}}+\underbrace{\left\langle{u_{j}^{*}\delta h}\right\rangle-\frac{1}{2}\left\langle{\omega_{j}^{*}\delta q^{2}}\right\rangle}_{\text{Turbulent transp.}}+\underbrace{\frac{1}{\rho}\left\langle{\delta p\delta u_{j}}\right\rangle}_{\text{Pressure transp.}}$ (7)
$\underbrace{-\frac{\nu}{2}\frac{\partial\left\langle{\delta h}\right\rangle}{\partial X_{j}}}_{\text{Viscous diff.}}\underbrace{-\frac{1}{2\rho}\left(\left\langle{\delta T^{\omega}_{ij}\delta u_{i}}\right\rangle+\left\langle{\delta T_{ij}\delta\omega_{i}}\right\rangle\right),}_{\text{Polymeric transp.}}$
where $j=1,2,3$ for the general case with three non homogeneous directions. The components of the flux vector in the space of scales $\bm{\phi}$ are:
$\phi_{j}^{\delta q^{2}}=\underbrace{\left\langle{\delta U_{j}\delta q^{2}}\right\rangle}_{\text{Mean transp.}}+\underbrace{\left\langle{\delta u_{j}\delta q^{2}}\right\rangle}_{\text{Turbulent transp.}}\underbrace{-2\nu\frac{\partial\left\langle{\delta q^{2}}\right\rangle}{\partial r_{j}}}_{\text{Viscous diff.}}\underbrace{-\frac{4}{\rho}\left\langle{T_{ij}^{*}\delta u_{i}}\right\rangle,}_{\text{Polymeric transp.}}$ (8)
$\phi_{j}^{\delta\omega^{2}}=\underbrace{\left\langle{\delta U_{j}\delta\omega^{2}}\right\rangle}_{\text{Mean transp.}}+\underbrace{\left\langle{\delta u_{j}\delta\omega^{2}}\right\rangle}_{\text{Turbulent transp.}}\underbrace{-2\nu\frac{\partial\left\langle{\delta\omega^{2}}\right\rangle}{\partial r_{j}}}_{\text{Viscous diff.}}\underbrace{-\frac{4}{\rho}\left\langle{T^{\omega*}_{ij}\delta\omega_{i}}\right\rangle,}_{\text{Polymeric transp.}}$ (9)
$\phi_{j}^{\delta h}=\underbrace{\left\langle{\delta U_{j}\delta h}\right\rangle}_{\text{Mean transp.}}+\underbrace{\left\langle{\delta u_{j}\delta h}\right\rangle-\frac{1}{2}\left\langle{\delta\omega_{j}\delta q^{2}}\right\rangle}_{\text{Turbulent transp.}}\underbrace{-2\nu\frac{\partial\left\langle{\delta h}\right\rangle}{\partial r_{j}}}_{\text{Viscous diff.}}\underbrace{-\frac{2}{\rho}\left\langle{T_{ij}^{\omega*}\delta u_{i}}\right\rangle-\frac{2}{\rho}\left\langle{T_{ij}^{*}\delta\omega_{i}}\right\rangle,}_{\text{Polymeric transp.}}$ (10)
where $j=1,2,3$. The first terms on the right of equation 4 are the production terms,
$P^{\delta q^{2}}=\underbrace{-2\left\langle{u_{j}^{*}\delta u_{i}}\right\rangle\delta\left(\frac{\partial U_{i}}{\partial x_{j}}\right)-2\left\langle{\delta u_{i}\delta u_{j}}\right\rangle\left(\frac{\partial U_{i}}{\partial x_{j}}\right)^{*},}_{\text{Mean Prod. (Grad. $U$)}}$ (11)
$\begin{gathered}P^{\delta\omega^{2}}=\underbrace{-2\left\langle{u_{j}^{*}\delta\omega_{i}}\right\rangle\delta\left(\frac{\partial\Omega_{i}}{\partial x_{j}}\right)-2\left\langle{\delta u_{j}\delta\omega_{i}}\right\rangle\left(\frac{\partial\Omega_{i}}{\partial x_{j}}\right)^{*}}_{\text{Mean Prod. (Grad.$\Omega$)}}+\\ \underbrace{2\left\langle{\omega_{j}^{*}\delta\omega_{i}}\right\rangle\delta\left(\frac{\partial U_{i}}{\partial x_{j}}\right)+2\left\langle{\delta\omega_{i}\delta\omega_{i}}\right\rangle\left(\frac{\partial U_{i}}{\partial x_{j}}\right)^{*}}_{\text{Mean Production (Grad. $U$)}}+\\ \underbrace{2\left\langle{\delta\omega_{i}\Omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)}\right\rangle+2\left\langle{\delta\omega_{i}\delta\Omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*}}\right\rangle}_{\text{Mean Vortex Stretching}}+\\ \underbrace{2\left\langle{\delta\omega_{i}\omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)}\right\rangle+2\left\langle{\delta\omega_{i}\delta\omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*}}\right\rangle}_{\text{Turbulent Vortex Stretching}},\end{gathered}$ (12)
$P^{\delta h}=$ $\underbrace{-\left\langle{u_{j}^{*}\delta u_{i}}\right\rangle\delta\left(\frac{\partial\Omega_{i}}{\partial x_{j}}\right)-\left\langle{\delta u_{j}\delta u_{i}}\right\rangle\left(\frac{\partial\Omega_{i}}{\partial x_{j}}\right)^{*}}_{\text{Mean Prod. (Grad. $\Omega$)}}+\underbrace{\left\langle{\delta u_{i}\Omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)}\right\rangle+\left\langle{\delta u_{i}\delta\Omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*}}\right\rangle}_{\text{Mean Vortex stretching}}+$
$\underbrace{-\left\langle{u_{j}^{*}\delta\omega_{i}}\right\rangle\delta\left(\frac{\partial U_{i}}{\partial x_{j}}\right)-\left\langle{\delta u_{j}\delta\omega_{i}}\right\rangle\left(\frac{\partial U_{i}}{\partial x_{j}}\right)^{*}+\left\langle{\omega_{j}^{*}\delta u_{i}}\right\rangle\delta\left(\frac{\partial U_{i}}{\partial x_{j}}\right)+\left\langle{\delta\omega_{j}\delta u_{i}}\right\rangle\left(\frac{\partial U_{i}}{\partial x_{j}}\right)^{*}}_{\text{Mean Prod. (Grad. $U$)}},$ (13)
and the fluid-polymer exchange term
$\begin{gathered}\Pi^{\delta q^{2}}=-\frac{4}{\rho}\underbrace{\left\langle{T_{ij}\frac{\partial u_{i}}{\partial x_{j}}}\right\rangle^{*}}_{\left\langle{\pi^{\delta q^{2}}}\right\rangle^{*}},\qquad\Pi^{\delta\omega^{2}}=-\frac{4}{\rho}\underbrace{\left\langle{T_{ij}^{\omega}\frac{\partial\omega_{i}}{\partial x_{j}}}\right\rangle^{*}}_{\left\langle{\pi^{\delta\omega^{2}}}\right\rangle^{*}}\qquad\text{and}\\ \Pi^{\delta h}=-\frac{4}{\rho}\underbrace{\left(\frac{1}{2}\left\langle{T_{ij}\frac{\partial\omega_{i}}{\partial x_{j}}}\right\rangle^{*}+\frac{1}{2}\left\langle{T_{ij}^{\omega}\frac{\partial u_{i}}{\partial x_{j}}}\right\rangle^{*}\right).}_{\left\langle{\pi^{\delta h}}\right\rangle^{*}}\end{gathered}$ (14)
The last terms on the right hand side correspond to the dissipation
@@
$\mathcal{E}^{\delta q^{2}}=-4\underbrace{\nu\left\langle{\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{i}}{\partial x_{j}}}\right\rangle^{*}}_{\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle^{*}},\ \mathcal{E}^{\delta\omega^{2}}=-4\underbrace{\nu\left\langle{\frac{\partial\omega_{i}}{\partial x_{j}}\frac{\partial\omega_{i}}{\partial x_{j}}}\right\rangle^{*}}_{\left\langle{\varepsilon_{f}^{\delta\omega^{2}}}\right\rangle^{*}},\ \text{and}\ \mathcal{E}^{\delta h}=-4\underbrace{\nu\left\langle{\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial\omega_{i}}{\partial x_{j}}}\right\rangle^{*}}_{\left\langle{\varepsilon_{f}^{\delta h}}\right\rangle^{*}},$ (15)
and to the external forcing
$F^{\delta q^{2}}=2\left\langle{\delta f_{i}\delta u_{i}}\right\rangle,\ F^{\delta\omega^{2}}=2\left\langle{\delta f^{\omega}_{i}\delta\omega_{i}}\right\rangle,\ \text{and}\ F^{\delta h}=\left\langle{\delta f^{\omega}_{i}\delta u_{i}}\right\rangle+\left\langle{\delta f_{i}\delta\omega_{i}}\right\rangle.$ (16)
In the previous equations, the asterisk superscript $\cdot^{*}$ denotes the average between the positions $\bm{X}\pm\bm{r}/2$, i.e. $a^{*}=(a(\bm{X}+\bm{r}/2)+a(\bm{X}-\bm{r}/2))/2$.

### v2

@@ 2.1 The governing equations
$\frac{\partial u_{i}}{\partial t}+u_{j}\frac{\partial u_{i}}{\partial x_{j}}=$ $\ -\frac{1}{\rho}\frac{\partial p}{\partial x_{i}}+\nu\frac{\partial^{2}u_{i}}{\partial x_{j}\partial x_{j}}+\frac{1}{\rho}\frac{\partial T_{ij}}{\partial x_{j}}+f_{i},$ (1)
$\frac{\partial R_{ij}}{\partial t}+u_{k}\frac{\partial R_{ij}}{\partial x_{k}}=$ $\ \frac{\partial u_{i}}{\partial x_{k}}R_{kj}+R_{ik}\frac{\partial u_{k}}{\partial x_{j}}-\frac{\left(\mathscr{P}R_{ij}-\delta_{ij}\right)}{\tau_{p}},$ (2)
$\frac{\partial u_{j}}{\partial x_{j}}=$ $\ 0,$ (3)
@@
$T_{ij}=\mu_{p}\frac{\left(\mathscr{P}R_{ij}-\delta_{ij}\right)}{\tau_{p}},$ (4)
where $\tau_{p}$ is the polymeric relaxation time, $\mu_{p}$ is the polymeric viscosity and $\delta_{ij}$ is the Kronecker delta. $\mathscr{P}$ is the Peterlin function and is equal to $\mathscr{P}=1$ for the purely elastic Oldroyd-B model and to $\mathscr{P}=(\ell_{max}^{2}-3)/(\ell_{max}^{2}-R_{ii})$ for the FENE-P model; $\ell_{max}$ is the maximum polymer extensibility, $R_{ii}$ is the instantaneous measure of the end-to-end length of the polymers.
@@
$\frac{\partial\omega_{i}}{\partial t}+u_{j}\frac{\partial\omega_{i}}{\partial x_{j}}=\omega_{j}\frac{\partial u_{i}}{\partial x_{j}}+\nu\frac{\partial^{2}\omega_{i}}{\partial x_{j}\partial x_{j}}+\frac{1}{\rho}\frac{\partial T_{ij}^{\omega}}{\partial x_{j}}+f_{i}^{\omega},$ (5)
@@
The flow is described by its mean and fluctuating fields, defined after Reynolds decomposition. The mean field is obtained by means of the $\left\langle{\cdot}\right\rangle$ operator that denotes averaging among realisations, along homogeneous directions, and in time if the flow is statistically stationary. Hereafter, capital letters ($U_{i}$, $\Omega_{i}$ and $P$) refer to mean quantities, while small letters ($u_{i}$, $\omega_{i}$ and $p$) to the fluctuations around them.
@@ 2.2 The budget equations
Following the work of Baj et al. (2022), we consider three specific structure functions, i.e. the velocity structure function $\delta q^{2}=\delta u_{i}\delta u_{i}$, the vorticity structure function $\delta\omega^{2}=\delta\omega_{i}\delta\omega_{i}$ and the helicity structure function $\delta h=\delta u_{i}\delta\omega_{i}$ where repeated indices imply summation. The three structure functions feature velocity and vorticity increments ($\delta u_{i}$ and $\delta\omega_{i}$) between two points $\bm{x}$ and $\bm{x}^{\prime}$, that can be identified by means of their midpoint $\bm{X}=(\bm{x}+\bm{x}^{\prime})/2$ and separation vector $\bm{r}=\bm{x}^{\prime}-\bm{x}$, namely
@@
The quantity $\left\langle{\delta q^{2}}\right\rangle(\bm{X},\bm{r},t)$ is commonly used as a proxy for the kinetic energy associated with eddies of scale $r=|\bm{r}|$ at position $\bm{X}$. This interpretation is however only approximate. Actually, $\left\langle{\delta q^{2}}\right\rangle$ reflects the cumulative energy contained in eddies with scales up to $r$, rather than the energy at that scale. This becomes clear when expressing $\left\langle{\delta q^{2}}\right\rangle$ in terms of the velocity variance and the correlation function $C_{uu}$. In the idealised case of HIT, we have $\left\langle{\delta q^{2}}\right\rangle=2\left\langle{u_{i}u_{i}}\right\rangle(1-C_{uu}(r))$, where $C_{uu}=\left\langle{u_{i}(\bm{x}+\bm{r},t)u_{i}(\bm{x},t)}\right\rangle/\left\langle{u_{i}u_{i}}\right\rangle$: $\left\langle{\delta q^{2}}\right\rangle=0$ for $r=0$ and $\left\langle{\delta q^{2}}\right\rangle\rightarrow 2\left\langle{u_{i}u_{i}}\right\rangle$ as $r\rightarrow\infty$. However, as pointed out by Davidson & Pearson (2005), this is not entirely accurate: $\left\langle{\delta q^{2}}\right\rangle$ at a given scale $r$ includes contributions not only from eddies of scale $\leq r$, but also from large-scale eddies via their associated enstrophy. A similar interpretation applies to the enstrophy and helicity structure functions. For example, for enstrophy in HIT we have $\left\langle{\delta\omega^{2}}\right\rangle=2\left\langle{\omega_{i}\omega_{i}}\right\rangle(1-C_{\omega\omega}(r))$, where $C_{\omega\omega}=\left\langle{\omega_{i}(\bm{x}+\bm{r},t)\omega_{i}(\bm{x},t)}\right\rangle/\left\langle{\omega_{i}\omega_{i}}\right\rangle\rightarrow 0$ and $\left\langle{\delta\omega^{2}}\right\rangle\rightarrow 2\left\langle{\omega_{i}\omega_{i}}\right\rangle$ when $r\rightarrow\infty$. The enstrophy structure function $\left\langle{\delta\omega^{2}}\right\rangle$ can be (loosely) interpreted as the cumulative enstrophy contributed by eddies of scales up to $r$.
@@
The budget equations for $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ describe production, transport and dissipation of energy, enstrophy and helicity in the space of scales $\bm{r}$ and positions $\bm{X}$. These equations link the variation in time of the three structure functions at a given scale and position to the instantaneous imbalance among production, transport and dissipation. Specifically, they correspond to transport budget equations governing the cumulative energy, enstrophy, and helicity associated with eddies of size up to scale $r$. The three budget equations are obtained after manipulation of the Navier–Stokes equations for the velocity and the vorticity, without any assumption of homogeneity and isotropy; for the full derivation (for a purely Newtonian fluid) of the equation for $\left\langle{\delta q^{2}}\right\rangle$ we refer the interested reader to the appendix of Gatti et al. (2020). The derivation of the budget equations for $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ starts with the evolution equations for velocity and vorticity, and requires a sequence of manipulations that use the incompressibility constraint. A brief recap of the main steps is also provided in Baj et al. (2022).
@@
In compact form, the budget equations for $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ can be written as:
@@
$\frac{\partial\left\langle{\delta g}\right\rangle}{\partial t}+\frac{\partial\langle\psi^{\delta g}_{j}\rangle}{\partial X_{j}}+\frac{\partial\langle\phi^{\delta g}_{j}\rangle}{\partial r_{j}}-\left\langle{\mathcal{P}^{\delta g}}\right\rangle-\left\langle{\mathfrak{p}^{\delta g}}\right\rangle-\langle\mathcal{D}_{f}^{\delta g}\rangle-\left\langle{\mathcal{F}^{\delta g}}\right\rangle=-\left\langle{\mathcal{E}^{\delta g}}\right\rangle-\left\langle{\Pi^{\delta g}}\right\rangle$ (6)
Here, $\langle\delta g\rangle$ denotes one of the three structure functions. The terms in the corresponding balance equation are defined as follows:
  • $\partial\langle\delta g\rangle/\partial t$ represents the temporal evolution of the structure function.
  • $\partial\langle\psi_{j}^{\delta g}\rangle/\partial X_{j}$ accounts for transport in physical space, arising from fluid–fluid and fluid–polymer interactions.
  • $\partial\langle\phi_{j}^{\delta g}\rangle/\partial r_{j}$ describes the scale-space transfer associated with these interactions.
  • $\langle\mathcal{P}^{\delta g}\rangle$ denotes the production term due to the interaction between the mean shear and the velocity fluctuations.
  • $\langle\mathfrak{p}^{\delta g}\rangle$ arises from the coupling between the strain-rate and the vorticity fluctuations. This term, associated with vortex stretching, appears only in the equation for $\langle\delta\omega^{2}\rangle$.
  • $\langle\mathcal{D}_{f}^{\delta g}\rangle$ represents viscous diffusion in both physical and scale spaces.
  • $\langle\mathcal{F}^{\delta g}\rangle$ is the contribution from the external forcing.
  • $\langle\mathcal{E}^{\delta g}\rangle$ denotes the two-point average of the fluid dissipation rate.
  • $\langle\Pi^{\delta g}\rangle$ denotes the two-point average of the polymeric source/sink term.
@@
We now provide the explicit form of terms appearing in the budget equations for the three structure functions. The components of the flux vector $\bm{\psi}$ in the physical space are given by:
$\psi_{j}^{\delta q^{2}}=\$$U_{j}^{*}\delta q^{2}+u_{j}^{*}\delta q^{2}+2\delta p\delta u_{j}/\rho-\delta T_{ij}\delta u_{i}/\rho,$ (7)
$\psi_{j}^{\delta\omega^{2}}=\$$U_{j}^{*}\delta\omega^{2}+u_{j}^{*}\delta\omega^{2}-\delta T_{ij}^{\omega}\delta\omega_{i}/\rho,$ (8)
$\psi_{j}^{\delta h}=\$$U_{j}^{*}\delta h+u_{j}^{*}\delta h-\omega_{j}^{*}\delta q^{2}/2+\delta p\delta u_{j}/\rho-(\delta T_{ij}^{\omega}\delta u_{i}+\delta T_{ij}\delta\omega_{i})/(2\rho)$ (9)
where $j=1,2,3$ correspond to the three spatial directions in the case of full inhomogeneity. The superscript ``*" denotes an arithmetic average between the two points $\bm{X}\pm\bm{r}/2$. The components of the flux vector $\bm{\phi}$ in the scale-space are given by:
$\phi_{j}^{\delta q^{2}}=\$$\delta U_{j}\delta q^{2}+\delta u_{j}\delta q^{2}-4T_{ij}^{*}\delta u_{i}/\rho,$ (10)
$\phi_{j}^{\delta\omega^{2}}=\$$\delta U_{j}\delta\omega^{2}+\delta u_{j}\delta\omega^{2}-4T_{ij}^{\omega,*}\delta\omega_{i}/\rho,$ (11)
$\phi_{j}^{\delta h}=\$$\delta U_{j}\delta h+\delta u_{j}\delta h-\delta\omega_{j}\delta q^{2}/2-2(T_{ij}^{\omega,*}\delta u_{i}+T_{ij}^{*}\delta\omega_{i})/\rho$ (12)
where $j=1,2,3$. The production terms $\mathcal{P}$ are given by:
$\mathcal{P}^{\delta q^{2}}=$ $\ -2u_{j}^{*}\delta u_{i}\delta\left(\frac{\partial U_{i}}{\partial x_{j}}\right)-2\delta u_{i}\delta u_{j}\left(\frac{\partial U_{i}}{\partial x_{j}}\right)^{*},$ (13)
$\mathcal{P}^{\delta\omega^{2}}=$ $\ -2u_{j}^{*}\delta\omega_{i}\delta\left(\frac{\partial\Omega_{i}}{\partial x_{j}}\right)-2\delta u_{j}\delta\omega_{i}\left(\frac{\partial\Omega_{i}}{\partial x_{j}}\right)^{*}+2\omega^{*}\delta\omega_{i}\delta\left(\frac{\partial U_{i}}{\partial x_{j}}\right)+2\delta\omega_{i}\delta\omega_{j}\left(\frac{\partial U_{i}}{\partial x_{j}}\right)^{*}+$
$\ +2\delta\omega_{i}\Omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)+2\delta\omega_{i}\delta\Omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*},$ (14)
$\mathcal{P}^{\delta h}=$ $\ -u_{j}^{*}\delta u_{i}\delta\left(\frac{\partial\Omega_{i}}{\partial x_{j}}\right)-\delta u_{i}\delta u_{j}\left(\frac{\partial\Omega_{i}}{\partial x_{j}}\right)^{*}+\delta u_{i}\Omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)+\delta u_{i}\delta\Omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*}+$
$\ -u_{j}^{*}\delta\omega_{i}\delta\left(\frac{\partial U_{i}}{\partial x_{j}}\right)-\delta u_{j}\delta\omega_{i}\left(\frac{\partial U_{i}}{\partial x_{j}}\right)^{*}+\omega_{j}^{*}\delta u_{i}\delta\left(\frac{\partial U_{i}}{\partial x_{j}}\right)+\delta\omega_{j}\delta u_{i}\left(\frac{\partial U_{i}}{\partial x_{j}}\right)^{*}.$ (15)
The additional generation term $\mathfrak{p}^{\delta\omega^{2}}$, appears only in the budget equation for $\langle\delta\omega^{2}\rangle$ and reads:
$\mathfrak{p}^{\delta\omega^{2}}=\ 2\delta\omega_{i}\omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)+2\delta\omega_{i}\delta\omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*}.$ (16)
The viscous terms $\mathcal{D}_{f}$, instead, are given by:
$\mathcal{D}_{f}^{\delta q^{2}}=$ $\ \frac{\nu}{2}\frac{\partial^{2}\delta q^{2}}{\partial X_{j}^{2}}+2\nu\frac{\partial^{2}\delta q^{2}}{\partial r_{j}^{2}},$ (17)
$\mathcal{D}_{f}^{\delta\omega^{2}}=$ $\ \frac{\nu}{2}\frac{\partial^{2}\delta\omega^{2}}{\partial X_{j}^{2}}+2\nu\frac{\partial^{2}\delta\omega^{2}}{\partial r_{j}^{2}},$ (18)
$\mathcal{D}_{f}^{\delta h}=$ $\ \frac{\nu}{2}\frac{\partial^{2}\delta h}{\partial X_{j}^{2}}+2\nu\frac{\partial^{2}\delta h}{\partial r_{j}^{2}},$ (19)
while the forcing terms $\mathcal{F}$ take the form:
$\mathcal{F}^{\delta q^{2}}=$ $\ 2\delta f_{i}\delta u_{i}$ (20)
$\mathcal{F}^{\delta\omega^{2}}=$ $\ 2\delta f_{i}^{\omega}\delta\omega_{i},$ (21)
$\mathcal{F}^{\delta h}=$ $\ \delta f_{i}^{\omega}\delta u_{i}+\delta f_{i}\delta\omega_{i}.$ (22)
The two terms at the right-hand side of equation (6) correspond to the two-point averages of the fluid dissipation and polymeric source/sink rates. The dissipation terms $\mathcal{E}$ are:
$\mathcal{E}^{\delta q^{2}}=$ $\ 4\varepsilon_{f}^{\delta q^{2},*}=4\nu\left(\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{i}}{\partial x_{j}}\right)^{*},$ (23)
@@
$\mathcal{E}^{\delta\omega^{2}}=$ $\ 4\varepsilon_{f}^{\delta\omega^{2},*}=4\nu\left(\frac{\partial\omega_{i}}{\partial x_{j}}\frac{\partial\omega_{i}}{\partial x_{j}}\right)^{*},$ (24)
$\mathcal{E}^{\delta h}=$ $\ 4\varepsilon_{f}^{\delta h,*}=4\nu\left(\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial\omega_{i}}{\partial x_{j}}\right)^{*}$ (25)
and the polymeric source/sink terms $\Pi$ are:
$\Pi^{\delta q^{2}}=$ $\ 4\pi^{\delta q^{2},*}/\rho=\frac{4}{\rho}\left(T_{ij}\frac{\partial u_{i}}{\partial x_{j}}\right)^{*},$ (26)
$\Pi^{\delta\omega^{2}}=$ $\ 4\pi^{\delta\omega^{2},*}/\rho=\frac{4}{\rho}\left(T_{ij}^{\omega}\frac{\partial\omega_{i}}{\partial x_{j}}\right)^{*},$ (27)
$\Pi^{\delta h}=$ $\ 4\pi^{\delta h,*}/\rho=\frac{4}{\rho}\left(\frac{1}{2}\left(T_{ij}\frac{\partial\omega_{i}}{\partial x_{j}}\right)+\frac{1}{2}\left(T_{ij}^{\omega}\frac{\partial u_{i}}{\partial x_{j}}\right)\right)^{*}.$ (28)
Equation (6) can be interpreted as a balance equation for the two-point average of the total dissipation rate. At each scale $\bm{r}$ and position $\bm{X}$, the sum of the terms on the left-hand side must balance the combined contribution of $\langle\mathcal{E}\rangle$ and $\langle\Pi\rangle$.